OpenStudy (anonymous):

two parallel resistors 6 Ohms each, \[\frac{ 1 }{ 6 } + \frac{ 1 }{ 6 } = \frac{ 1 }{ 12 }\] what is total resistance?

hero (hero):

If I remember correctly, it should be this: \[\frac{6 \times 6}{6 + 6}\]

OpenStudy (anonymous):

looks good

hero (hero):

And in general the formula for total resistance it: \[R_{\text{total}} = \frac{R_1 \times R_2}{R_1 + R_2}\]